OpenStudy (anonymous):

A car sits in an entrance ramp to a freeway, waiting for a break in the traffic. The driver sees a small gap between a van and an 18-wheel truck and accelerates with constant acceleration along the ramp and onto the freeway. The car starts from rest, moves in a straight line, and has a speed of 25.0 when it reaches the end of the ramp, which has length 129 .

OpenStudy (anonymous):

What is the acceleration of the car?s

OpenStudy (anonymous):

\[(V _{f})^{2} = (V _{o})^{2} + 2ad\] (25)^2 = o^2 + 2(129)a 50 = 258a a = 0.194 distance/time/time

OpenStudy (anonymous):

I realized just now that I gave you an incorrect answer. The setup was correct, but I made a silly error in the math. 25^2 is 625, not 50. 625/258 is 2.42, so the correct answer should be 2.42 distance/time/time